OpenStudy (anonymous):

i need help with a problem is any1 a teacher or professor of any sort ?? a medal is awarded if you can give me the correct answer.

OpenStudy (ddcamp):

What's your question?

OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. X(0, 6), Y(5, 6)

OpenStudy (ddcamp):

The Standard Form of an equation for a line looks like: \[Ax + By = C\] To find this, start by finding the equation in point-slope form: \[y-y_1=m(x-x_1) \\ (x_1,y_1)\ is\ a\ point\ on\ the\ line \\ m\ is\ the\ slope\ of\ the\ line\]

OpenStudy (anonymous):

thanx so much your a life saver

OpenStudy (ddcamp):

To find m, you can use the equation: \[m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

0/-5 zero slope rite

OpenStudy (ddcamp):

Yes, which just equals 0

OpenStudy (anonymous):

ok so y-y1=0

OpenStudy (ddcamp):

No, y-y1 = y-6 If the y doesn't have a subscript, its a variable.